Find the distance between (6, –2) and (–6, 3).
Answer by checkley75(3666)

THE X SIDE IS (6+6)=12
THE Y SIDE IS (3+2)=5
SO THE DISTANCE IS THE HYPOTENUSE THUS
12^2+5^2=H^2
144+25=H^2
169=H^2
H=SQRT169
H=13 IS THE HYPOTENUSE & THE DISTANCE BETWEEN THESE POINTS.